# Deconfinement and chiral phase transitions in quark matter with chiral imbalance

### v1

@@ IV Phase diagram at finite quark chemical potential
Finally, we consider the influence on the phase diagram in the $T\times\mu$ plane of including the dependence on $\mu_{5}$ in the Polyakov loop potential as indicated in Eq. (37). In this case, we have $\Phi\neq\Phi^{\dagger}$ and, instead of Eq. (35), we need to solve
@@ V Conclusions
In this paper, we performed a thorough study within the PNJL model on the differences that the TRS and MSS regularization procedures imply for the thermodynamics of quark matter with chiral imbalance. We have shown that, to reproduce lattice results for the chiral and deconfinement pseudocritical temperatures as a function of the chiral chemical potential $\mu_{5}$, it is necessary to include a $\mu_{5}$-dependece in the parametrization of the quadratic $\Phi^{\dagger}\Phi$ term in Polyakov loop potential, in addition to the usual temperature dependence. We have proposed a polynomial dependence analogous to that in terms of temperature. In addition to reproducing the lattice results for values of the pseudocritical temperatures, the proposed parametrization also reproduces the increase of those temperatures with $\mu_{5}$, indicating that such a $\mu_{5}$ dependence in the Polyakov loop potential may be the missing ingredient in the traditional PNJL models. We have studied the changes caused by this new parametrization of the Polyakov loop potential in various thermodynamic quantities, and also contrasted the results predicted by the TRS and MSS regularization procedures. In particular, we have shown that the combined use of the new parametrization of the Polyakov loop potential and the MSS procedure leads to consistent results that agree with those of lattice simulations.
VI Acknowledgements
This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grants No. 312032/2023-4 (R.L.S.F.), 307286/2021-5 (R.O.R.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants Nos. 19/2551-0000690-0 and 19/2551-0001948-3 (R.L.S.F.), and 23/2551-0000791-6 and 23/2551-0001591-9 (D.C.D.); Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021 (R.O.R.); Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 (F.X.A.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia - Física Nuclear e Aplicações (INCT - FNA), Grant No. 464898/2014-5 and supported by the Serrapilheira Institute (grant number Serra - 2211-42230).

### v2

@@ IV Phase diagram at finite quark chemical potential
Finally, we consider the influence on the phase diagram in the $T\times\mu$ plane of including the dependence on $\mu_{5}$ in the Polyakov-loop potential as indicated in Eq. (37). In this case, we have $\Phi\neq\Phi^{\dagger}$ and, instead of Eq. (35), we need to solve
@@ V Conclusions
In this paper, we performed a thorough study within the PNJL model on the differences that the TRS and MSS regularization procedures imply for the thermodynamics of quark matter with chiral imbalance. We have shown that, to reproduce lattice results for the chiral and deconfinement pseudocritical temperatures as a function of the chiral chemical potential $\mu_{5}$, it is necessary to include a $\mu_{5}$-dependece in the parametrization of the quadratic $\Phi^{\dagger}\Phi$ term in Polyakov-loop potential, in addition to the usual temperature dependence. We have proposed a polynomial dependence analogous to that in terms of temperature. In addition to reproducing the lattice results for values of the pseudocritical temperatures, the proposed parametrization also reproduces the increase of those temperatures with $\mu_{5}$, indicating that such a $\mu_{5}$ dependence in the Polyakov-loop potential may be the missing ingredient in the traditional PNJL models. We have studied the changes caused by this new parametrization of the Polyakov-loop potential in various thermodynamic quantities, and also contrasted the results predicted by the TRS and MSS regularization procedures. In particular, we have shown that the combined use of the new parametrization of the Polyakov-loop potential and the MSS procedure leads to consistent results that agree with those of lattice simulations.
Acknowledgements
This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grants No. 312032/2023-4 (R.L.S.F.), No. 307286/2021-5 (R.O.R.) and No. 309262/2019-4 (G.K.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants No. 19/2551-0000690-0 and 19/2551-0001948-3 (R.L.S.F.), and 23/2551-0000791-6 and 23/2551-0001591-9 (D.C.D.); Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) Grant No. 2018/25225-9 (G.K.); Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021 (R.O.R.); Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 (F.X.A.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia - Física Nuclear e Aplicações (INCT - FNA), Grant No. 464898/2014-5 and supported by the Serrapilheira Institute (Grant No. Serra - 2211-42230).